OpenStudy (anonymous):

Circle S has an equation of (x – 16)2 + (y + 9)2 = 4. What is the center and radius of circle S? Center: (16, -9); Radius: 4 Center: (-16, 9); Radius: 2 Center: (16, -9); Radius: 2 Center: (-16, 9); Radius: 4

OpenStudy (anonymous):

write the equation in general form

OpenStudy (anonymous):

is it b?

OpenStudy (anonymous):

it is 1...because u must change the signs

OpenStudy (anonymous):

it is the third one

OpenStudy (anonymous):

The general equation of a circle is (x-a)^2 +(y-b)^2 = r^2 Center is (a,b) Radius = r

OpenStudy (anonymous):

oh,yeah @beketso is correct

OpenStudy (anonymous):

3

OpenStudy (anonymous):

From the equation you have: (x-16)^2+(y-(-9))^2=2^2 the correct answer is : center: (16, -9) radius: 2

OpenStudy (anonymous):

so is it 1 or 3 O.o

OpenStudy (anonymous):

thank you guyss sooo muchhh !

OpenStudy (anonymous):

pleasure